### v2

# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

Abstract
While auxiliary information has become a key to enhance Large Language Models (LLMs), relatively little is known about how LLMs merge these contexts, specifically generated and retrieved. To study this, we formulate a systematic framework to identify whether LLMs’ responses, derived from the integration of generated and retrieved contexts, are attributed to either generated or retrieved contexts. To achieve this, we construct datasets with conflicting contexts, where each question is paired with both generated and retrieved contexts, yet only one of them contains the correct answer. Our experiments reveal a significant bias in LLMs (GPT-4/3.5 and Llama2) towards generated contexts, even when they provide incorrect information. We further identify two key factors contributing to this bias: i) contexts generated by LLMs typically show greater similarity to the questions, increasing their likelihood of selection; ii) the segmentation process used in retrieved contexts disrupts their completeness, thereby hindering their full utilization in LLMs. Our analysis enhances the understanding of how LLMs merge diverse contexts, offering valuable insights for advancing current augmentation methods for LLMs.
[…]
Blinded by Generated Contexts: How Language Models
Merge Generated and Retrieved Contexts for Open-Domain QA?
Hexiang ${}^{\spadesuit\heartsuit}$, Fei ${}^{\spadesuit}$22footnotemark: 2  , Wanli ${}^{\spadesuit\diamondsuit}$, Yuanzhuo ${}^{\spadesuit}$, Qi ${}^{\spadesuit}$, Xueqi ${}^{\spadesuit\heartsuit}$$^{\spadesuit}$ Key Laboratory of AI Safety & Security, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China ${}^{\heartsuit}$niversity of Chinese Academy of Sciences, Beijing, China ${}^{\diamondsuit}$ankai University, Tianjin, China {tanhexiang21s, sunfei, wangyuanzhuo, caoqi, cxq}@ict.ac.cn
1 Introduction
Recent advancements in augmenting Large Language Models (LLMs) with auxiliary information have significantly revolutionized their efficacy in knowledge-intensive tasks Chang et al. (2023); Ram et al. (2023). In this evolving field, existing works can be broadly categorized into two groups based on information sources: generation-augmented and retrieval-augmented approaches. To effectively harness the internal knowledge of LLMs, generation-augmented approaches Liu et al. (2022); Sun et al. (2023), e.g., GenRead Yu et al. (2022), instruct LLMs to initially generate a background context tailored to the given question, which is then employed as the basis for producing the final answer. In contrast, retrieval-augmented approaches Lewis et al. (2020); Ram et al. (2023) adopt an alternative strategy by incorporating relevant passages from external corpora, e.g., Wikipedia, as context, thereby notably enhancing LLMs’ capability to address challenges like knowledge updates Jang et al. (2022) and long-tail knowledge Kandpal et al. (2023).
Figure 1: Blue bars show the success number on NQ test set with only retrieved contexts, while warm orange bars depict the decline in success for the same questions when generated contexts are additionally incorporated.
Building on the foundations laid by generation-augmented and retrieval-augmented methods, recent hybrid approaches have attempted to integrate them to further improve performance in tasks such as Question Answering (QA) Yu et al. (2022); Mallen et al. (2023). These hybrid approaches face a significant challenge: conflicts between diverse sources can impede the effectiveness of information integration Zhang et al. (2023). While recent works have investigated conflicts within contexts from a single source (Xie et al., 2023; Chen et al., 2022), it remains unclear how LLMs resolve conflicts between generated and retrieved contexts. This study, therefore, aims to investigate the underlying mechanisms by which LLMs process the two types of contexts, especially when they contain conflicting information.
Our investigation was driven by a striking observation: in certain cases, models relying solely on retrieval contexts succeeded, whereas, counter-intuitively, hybrid approaches failed, as depicted in Figure 1. To uncover the underlying reasons, we proposed a systematic framework to dissect the process by which LLMs merge generated and retrieved contexts. We curated tailored datasets in which each question is accompanied by a pair of generated and retrieved contexts. These contexts are intentionally designed to be inconsistent with each other, and only one of them contains the correct answer to the corresponding question. These datasets provide an innovative lens to unravel how LLMs discern and integrate diverse contextual information in QA.
In this paper, we conducted a series of controlled experiments using our uniquely designed datasets to empirically study this question, focusing on several state-of-the-art closed (GPT-3.5/4) and open (Llama2-7b/13b) LLMs. Surprisingly, our findings reveal a pronounced bias in LLMs to favor generated contexts, even when the generated contexts offer incorrect information while the retrieval contexts hold the correct answers. Furthermore, this bias persists regardless of whether the generated text was produced internally or by other LLMs. These findings highlight a critical challenge for existing LLMs in effectively merging internal parametric knowledge (i.e., generated contexts) and external information (i.e., retrieved contexts), under increasingly common non-tunable settings, e.g., those involving black-box APIs like GPT-4.
To investigate the underlying causes of the bias, we study the role of confirmation bias Xie et al. (2023) (Section 5.1), text similarity (Section 5.2) and semantic completeness (Section 5.3). We find that: i) LLMs maintain a significant preference for generated contexts when they contain information inconsistent with LLMs’ parametric knowledge. ii) Generated contexts typically exhibit a higher degree of similarity to the questions compared to retrieved contexts, even when they contain incorrect information. This similarity gap is a significant factor in LLMs’ preference for generated contexts, which emphasizes the the need for caution with LLM-generated contexts, to avoid being misled by highly relevant but inaccurate information. iii) LLMs exhibit significant sensitivity to the semantic completeness of input contexts. The segmentation process used in retrieved contexts may disrupt their completeness, thereby hindering their full utilization in LLMs.
This work preliminarily explores an growing challenge for LLMs to utilize contexts from diverse sources, especially as content generated by LLMs, which may include potential misinformation (Pan et al., 2023), becomes more prevalent on the internet. Furthermore, our findings offer valuable guidance for enhancing existing retrieval-augmented methods, such as optimizing passage segmentation in retrieval systems. Our main contributions can be summarized as:
  • We uncover a critical bias in existing LLMs, where they heavily rely on generated contexts regardless of correctness, indicating an insufficient use of diverse information sources.
  • To facilitate controlled experiments, we develop a specialized framework for constructing tailored datasets and excluding confounding factors, e.g., input order and context length.
  • Our extensive analyses have identified two key factors, i.e., text similarity and semantic completeness, in the context utilization of LLMs. Moreover, the recently identified confirmation bias (Xie et al., 2023) plays a relatively minor role.
2 Background & Study Formulation
(a) Retrieval-Augmented Approach
(b) Generation-Augmented Approach
(a) Retrieval-Augmented Approach
(b) Generation-Augmented Approach
(c) Hybrid Approach
Figure 2: The frameworks of retrieval-augmented approach, generation-augmented approach and hybrid approach.
In this section, we briefly review three categories of LLMs augmented with auxiliary information for QA tasks: retrieval-augmented, generation-augmented, and hybrid approaches. Additionally, we introduce the framework of our investigation.
2.1 Background
Figure 2 presents high-level abstract frameworks for three typical types of QA systems, each centered around an LLM as the reader component, and potentially incorporating additional components like a retriever, generator, or a blend of both, tailored to the specific methodology.
Retrieval-Augmented Approach. As shown in Figure 2, for a given question $q$ in a set of questions $\mathbb{Q}$, these approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) initially use a retrieval model $\gamma$ to select the top $k$ relevant documents $D_{k}^{\gamma}{=}\gamma_{k}(q,\mathbb{C}){=}\{d_{1}^{\gamma},\dots,d_{k}^{\gamma}\}$ from a corpus =$\mathbb{C}{=}\{d_{1},\dots,d_{|\mathbb{C}|}\}$. Then, a reader (often LLM) $\phi$ employs these documents $D^{\gamma}$ to generate an answer $a_{\phi}^{\gamma}$, expressed as $a_{\phi}^{\gamma}=\phi(q,D^{\gamma})$.
Generation-Augmented Approach. In contrast, as illustrated in Figure 2, these works (Yu et al., 2022; Sun et al., 2023; Liu et al., 2022) involve an LLM as a generator $\varrho$ to produce $k$ tailored background contexts $D_{k}^{\varrho}{=}\varrho_{k}(q){=}\{d_{1}^{\varrho},\dots,d_{k}^{\varrho}\}$ for a give question $q$, thereby enhancing the utilization of the LLM’s internal knowledge. These LLM-generated contexts $D^{\varrho}$ form the input for reader $\phi$ to produce the final answer: $a_{\phi}^{\varrho}=\phi(q,D^{\varrho})$.
Hybrid Approach, as depicted in Figure 2, combines retrieved and generated contexts to enhance performance Yu et al. (2022); Abdallah and Jatowt (2023), as $a_{\phi}=\phi(q,D_{k}^{\gamma},D_{k}^{\varrho})$. These hybrid approaches face a significant challenge: conflicts between diverse sources can impede the effectiveness of information integration Zhang et al. (2023).
Knowledge Conflicts within Contexts. These studies mainly focus on conflicts within a single type of input contexts, either only retrieved (Chen et al., 2022) or generated (Xie et al., 2023), leaving underexplored how LLMs resolve conflicts between diverse contexts. A systematic review of related works is provided in Section A.1.
2.2 Answer Tracing Task
Figure 3: The task to study LLMs’ merging mechanisms by tracing the sources of the answers.
Departing from previous research, our study investigates the mechanisms by which LLMs merge contexts from diverse sources, i.e., hybrid approach setting. As illustrated in Figure 3, we design a task to ascertain whether an answer $a_{\phi}$ originates from generated contexts $D_{k}^{\varrho}$ or retrieved contexts $D_{k}^{\gamma}$. Our task mirrors the basic process of hybrid approaches,while our focus is studying LLMs’ merging mechanisms by tracing the sources of the answers instead of improving QA performance. For a more controlled and simpler analysis, we limit the context to a single instance from each source, i.e., $k{=}1$ and $a_{\phi}{=}\phi(q,d^{\gamma}_{1},d_{1}^{\varrho})$. Subsequently, by comparing the answer $a_{\phi}$ with the answers derived from the retrieved context $a_{\phi}^{\gamma}$ and the generated context $a_{\phi}^{\varrho}$, we can determine its source, thereby analyzing the merging mechanism of LLMs.
We specifically focus on non-tunable, i.e. zero-shot setting, LLMs acting as the reader and generator, reflecting prevalent real-world use cases like ChatGPT. This direction is motivated by the high cost and limited accessibility of fine-tuning, which makes the direct use of non-tunable LLMs popular. Additionally, given the extensive use of LLMs, any bias or issue in their merging mechanisms could lead to serious consequences.
3 Experimental Setup
To facilitate our investigation into how LLMs merge generated and retrieved contexts, this section elaborates on the construction of our context-conflicting datasets and the evaluation metric.
3.1 Context-Conflicting Datasets
Figure 4: The framework of constructing context-conflicting datasets .
As depicted in Figure 4, in our dataset $\mathcal{D}_{\mathrm{cc}}$, each sample $x$ is a quintet $(q,d_{1}^{\gamma},d_{1}^{\varrho},a_{\phi}^{\gamma},a_{\phi}^{\varrho})$, where $d_{1}^{\gamma}$ is the context returned by retriever $\gamma$ for question $q$, $d_{1}^{\varrho}$ represents the context generated by LLM $\varrho$, $a_{\phi}^{\gamma}$ and $a_{\phi}^{\varrho}$ are the candidate answers provided by the reader $\phi$, each based solely on the respective contexts $d_{1}^{\gamma}$ and $d_{1}^{\varrho}$. To guarantee that our dataset is suitable for controlled experiments aimed at investigating the merging mechanisms of LLMs, it must adhere to specific criteria:
  • [itemsep=3pt, topsep=0pt, partopsep=0pt, parsep=2pt, leftmargin=10pt]
  • Traceability: $a_{\phi}^{\gamma}$ and $a_{\phi}^{\varrho}$ must be supported by their respective contexts, $d_{1}^{\gamma}$ and $d_{1}^{\varrho}$.
  • Exclusivity: Only one of the contexts, $d_{1}^{\gamma}$ or $d_{1}^{\varrho}$ provides the correct answer, i.e., either $a_{\phi}^{\gamma}$ or $a_{\phi}^{\varrho}$ matches the gold answer of question $q$.
Such constraints establish a solid basis to identify which context, generated or retrieved, is selected by LLMs to produce answers in hybrid approaches.
We utilize the dev and test sets of two open-domain QA benchmark datasets with golden answers, i.e., NaturalQuestions (NQ) Kwiatkowski et al. (2019) and TriviaQA (TQA) Joshi et al. (2017), to assemble our experimental datasets. The overall pipeline for dataset construction is depicted in Figure 4, with detailed steps outlined as follows:
Context Preparation. Step 1 in Figure 4 illustrates the process of preparing retrieved and generated contexts for each question. For retrieved context, it is obtained from the top-$1$ ranked passage from Wikipedia using a retrieval model, e.g., Contriever (Izacard et al., 2021), a powerful off-the-shelf retrieval model that is extensively employed in various retrieval-augmented generation systems Shi et al. (2023); Ram et al. (2023).
For generated context, we follow the GenRead Yu et al. (2022), instructing the generator, i.e., LLM like GPT-4, to generate a background document based on the question. However, this method often yields contexts much longer (${>}250$ words) than the retrieved contexts (typically truncated to ${\sim}100$ words (Karpukhin et al., 2020; Izacard et al., 2021)). The discrepancy in length could potentially affect the merging mechanisms of LLMs, as discussed in Xie et al. (2023). To exclude this disturbance, we regulate the length of the generated context by incorporating length constraint in the prompt, resulting in an average length discrepancy below $3$%. All subsequent experiments, unless otherwise specified, employ this method to eliminate the impact of length variations. More details can be found in Appendix A.5.
Sample Filtering for Traceability. With each question paired with a single context (either generated or retrieved) established in the initial stage, the reader generates the corresponding candidate answer, as shown in Step 2 of Figure 4. To unravel the mechanisms of LLMs in context merging, it is essential to ensure the traceability, i.e., the output answer is derived from the input context, rather than the intrinsic parametric knowledge of the LLMs. To achieve this, we only keep samples in which both the generated and retrieved contexts exactly include their respective generated answers, exemplified by $a_{\phi}^{\gamma}{\in}d_{1}^{\gamma}$, where $\in$ denotes $d_{1}^{\gamma}$ contains the substring $a_{\phi}^{\gamma}$. This practice is grounded in the findings of Chen et al. (2022); Xie et al. (2023), which demonstrate that in the presence of external context, LLMs tend to rely on external context rather than their intrinsic parametric knowledge.
Building Context-Conflicting Dataset. Having obtained answers for each type of context, we are now positioned to construct our context-conflicting (CC) datasets, as depicted in Step 3 of Figure 4. Initially, We employ the exact match metric Yu et al. (2022) to evaluate the correctness of candidate answers derived from contexts, considering an answer correct if its normalized form matches any of the golden answers. Subsequently, the conflicting context datasets are composed of samples for which only one of the two types of contexts, either generated or retrieved, yields the correct answer, thereby ensuring the exclusivity. Notably, each dataset comprises two distinct subsets: AIG, consisting of samples with correct answers only in the generated context; and AIR comprising samples with correct answers only in the retrieved context.
3.2 Statistics of Datasets
For each reader-generator pair, we respectively constructs context-conflicting datasets from the original NQ and TQA test sets: NQ-CC (NQ-AIG + NQ-AIR) and TQA-CC (TQA-AIG + TQA-AIR).
We initially adopt a typical and simple setting in which a singular LLM serves as both the generator and reader. Table 1 provides detailed statistics for the constructed subsets corresponding to various LLMs, including GPT-4 (gpt-4-0613), GPT-3.5 (gpt-3.5-turbo-0613), Llama2-7b/13b (Llama-2-7b/13b-chat (Touvron et al., 2023)). The statistics show that the context-conflicting datasets form a substantial part of the test sets, underscoring the need to investigate how LLMs integrate these distinct contexts, as many questions are answerable with only one type. Notably, GPT-4 has fewer conflicting instances than other LLMs, because of its higher efficacy in answering questions using either solely retrieved or generated contexts.
Section 4.2 also explores a more complex scenario in which the generator and reader are distinct LLMs and show the statistics in Section A.7.
Generator &Reader NQ (12367) TQA (20150) NQ-AIG NQ-AIR TQA-AIG TQA-AIR GPT-4 1120 763 1712 681 GPT-3.5 1337 857 2389 1042 Llama2-13b 1441 1336 2982 2091 Llama2-7b 1423 1381 3064 2604 Avg. Prop. 10.8% 8.8% 12.6% 8.0%
Table 1: Dataset statistics across LLMs, “Avg. Prop.” shows average proportions of subsets to original datasets.
3.3 Evaluation Metric
Besides datasets, we also develop metrics to study how LLMs merge generated and retrieved contexts in hybrid approaches. Specifically, the selection of LLMs towards either generated or retrieved context can be measured by the proportion of answers that exactly match the answer produced solely by the corresponding context, denoted as
$\rho_{\text{gen}}{=}\text{avg}(\mathtt{em}(a_{\phi},a_{\phi}^{\varrho})),\quad\rho_{\text{ret}}=\text{avg}(\mathtt{em}(a_{\phi},a_{\phi}^{\gamma}))$
where $\mathtt{em}(a,b)$ returns 1 if $a$ exactly match $b$, and 0 otherwise. To facilitate a simple and efficient experiment, we define a synthesized metric as follows:
$\mathrm{DiffGR}=\frac{\rho_{\text{gen}}-\rho_{\text{ret}}}{\rho_{\text{gen}}+\rho_{\text{ret}}}$ (1)
The metric $\mathrm{DiffGR}$, ranging from $[-1,1]$, quantifies the extent of LLMs’ preference for choosing generated contexts over retrieved contexts. Using AIR as an example, where all correct answers come from retrieved contexts, an ideal $\mathrm{DiffGR}$ value should be $-1$, and $\mathrm{DiffGR}{>}0$ would indicate a significant bias towards generated contexts.
4 How LLMs Merge Contexts?
In this section, we conduct experiments on the proposed context-conflicting datasets to investigate the merging mechanism of the LLMs in hybrid approaches in different settings. We first consider a typical setting where generator and reader share a single LLM, to explore how LLMs merge retrieved and self-generated contexts (Section 4.2). Then, we extend our experiments to more flexible combinations of generator and reader (Section 4.2) to investigate their effects.
4.1 LLMs Prefer Self-Generated Contexts
Generator &Reader NQ-CC TQA-CC NQ-AIG NQ-AIR TQA-AIG TQA-AIR GPT-4 91.34 17.69 94.57 19.09 GPT-3.5 91.85 14.94 94.14 18.52 Llama2-13b 90.22 18.64 92.12 20.28 Llama2-7b 70.77 21.51 81.17 22.16
Table 2: The Exact Match (EM) scores (%) of hybrid approaches on corresponding context-conflicting datasets.
Figure 5: The $\mathrm{DiffGR}$ of LLMs on their corresponding context-conflicting datasets.
Our preliminary experiments, in which a single LLM serves both as generator and reader, are designed to explore how LLMs integrate information from retrieved and self-generated contexts. The LLMs under evaluation are tasked with answering questions using both types of contexts on their corresponding context-conflicting subsets. In all experiments, we employ a randomized input sequence of contexts to mitigate the influence of order, which is further discussed in Appendix A.8.
We begin our analysis by examining LLMs’ QA performance on context-conflicting datasets to reveal how well can LLMs utilize both types of contexts. Table 2 presents the Exact Match scores (Yu et al., 2022) across various LLMs. Surprisingly, LLMs demonstrate significantly low performance ($\leq 22.16\%$) on AIR subsets, despite the fact that the retrieved context alone consistently yields the correct answer on these subsets. In contrast, most LLMs exhibit strong performance on AIG subsets (near $90\%$ or more). An exception is the Llama2-7B model, which performs relatively poorly, likely due to its inferior instruction-following capabilities. Overall, all LLMs exhibit a significant performance gap between AIR and AIG datasets, with a pronounced decline in performance when the correct answers come from retrieved contexts.
To further reveal LLMs’ behavior underlying the QA performance, we trace the source contexts of LLMs’ answers using the proposed $\mathrm{DiffGR}$ metric. An ideal LLM that optimally leverages information from both sides should always rely on retrieved contexts in case of AIR (with $\mathrm{DiffGR=-1}$), and on generated contexts for AIG (with $\mathrm{DiffGR=1}$). Contrary to expectations, Figure 5 illustrates that LLMs fail to identify the correct information and consistently tend to rely on generated contexts on both AIG and AIR subsets. This result indicates a pronounced bias in LLMs to favor generated contexts, even when they provide incorrect information. This bias leads to the insufficient utilization of retrieved contexts mentioned above, and highlights a critical challenge for existing LLMs in effectively merging generated and retrieved contexts.
As the bias on AIR subsets has more direct impact on the performance, the following experiments and analysis will focus on the biases on these subsets to conserve space. Results on the AIG subsets can be found in Appendix A.9.
4.2 LLMs Broadly Prefer Generated Contexts
Figure 6: $\mathrm{DiffGR}$ with different (reader, generator) pairs on their corresponding NQ-AIR datasets.
Above experiments reveal the bias in LLMs to favor the self-generated context. Consequently, an intriguing question emerges: Are LLMs biased exclusively towards self-generated contexts, or do they exhibit a similar preference for contexts generated by other LLMs? To investigate this question, this section extends the experiments to more flexible combinations of generators and readers. This setting is also of practical significance, as recent works has explored the decoupling of generators and readers to achieve modularization of knowledge (Luo et al., 2023; Feng et al., 2023).
We construct context-conflicting datasets for each (generator, reader) pair respectively, as detailed in Appendix A.7. Base on these datasets, we then compute $\mathrm{DiffGR}$ metric to examine biases across various (generator, reader) pairs, as shown in Figure 6, and observe two notable insights: (i) LLMs also biased towards contexts generated by other LLMs.This suggests that such bias in LLMs is widespread and not limited to self-generated contexts. (ii) LLMs exhibit a stronger preference for contexts generated by themselves.The sole exception is Llama2-7b, which shows the strongest preference when paired with Llama2-13b as its generator. This phenomenon likely results from their highly similar model structures and training processes (Touvron et al., 2023).
5 Why LLMs Prefer Generated Contexts
In this section, we investigate why LLMs prefer generated contexts rather than retrieved contexts, from several perspectives: confirmation bias in Section 5.1, context similarity to the question in Section 5.2, and context completeness in Section 5.3.
5.1 Effect of Confirmation Bias
Recently, Xie et al. (2023) identified a confirmation bias in LLMs, wherein they exhibit a preference for contexts consistent with their parametric knowledge (or parametric memory), when faced with two conflicting generated contexts. In our experiments discussed in Section 4.1, which involve a single LLM serving both as reader and generator, it is natural to assume that the generated contexts align with the LLM’s parametric knowledge. This assumption leads to a key question: Does the confirmation bias lead to the observed preference for generated contexts?
To investigate the effect of confirmation bias, we designed controlled experiments that disrupt the consistency between generated contexts and the LLMs’ parametric knowledge. Overall, we enforce LLMs to make up a counter-memory context $d^{\varrho^{\prime}}_{1}$, which supports a same-type yet different answer compared to the original generated context. Then, we replace the generated context $d^{\varrho}_{1}$ with the counter-memory context $d^{\varrho^{\prime}}_{1}$ and recalculate $\mathrm{DiffGR}$ to assess shifts in preference after excluding confirmation bias. Table 3 reveals that LLMs maintain a significant preference for generated contexts when they contain information inconsistent with parametric knowledge. This indicates that confirmation bias does not play a major role in the preferences for generated contexts. Instead, other distinct differences between generated and retrieved contexts may be more influential, potentially overshadowing the effects of confirmation bias. In Section A.10, we further discuss the relationship between our conclusions and confirmation bias, and provide a detailed description of the experimental design process.
Context pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Gen vs. Ret 0.7561 0.6747 0.7058 0.6575 Ctr   vs. Ret 0.7342 0.6468 0.8010 0.6596
Table 3: $\mathrm{DiffGR}$ of different input context pairs. “Gen”, “Ret” and “Ctr” respectively represent generated contexts, retrieved contexts and counter-memory contexts.
5.2 Effect of Text Similarity
The text similarity between a context and a question can reflect the degree of their relevance. To investigate the potentially effect of the similarity, we employ Jaccard similarity and BERTScore (Zhang et al., 2020) to analyze the contexts on the constructed context-conflicting datasets with the reader and generator sharing a single LLM. Figure 7 shows that generated contexts exhibit a significant higher similarity to question, even on AIR subsets where generated contexts are incorrect. This similarity discrepancy persists whether assessed by term-based overlap ($0.37$ vs. $0.18$ on TQA-AIR) or semantic similarity ($0.90$ vs. $0.86$).
Figure 7: Context-question similarity distribution of generated and retrieved contextzs on the union of AIR subsets for different LLMs (GPT-4, GPT-3.5, Llama2-13b, Llama2-7b). Distribution on AIG subsets is presented in Appendix A.11, with similar results.
Figure 8: The $\mathrm{DiffGR}$ in slices with different average $\Delta\operatorname{sim}$, with results for other LLMs in Appendix A.11.
To further clarify the influence of the observed similarity discrepancies, we rank the samples according to the similarity gap $\Delta\operatorname{sim}$ between generated and retrieved contexts, and divide the dataset into $n$ ($n=5$ here [^1] Similar results and observations are found with other $n$.) slices with an equal number of samples.
$\Delta\operatorname{sim}=\frac{\operatorname{sim}(q,d^{\varrho})-\operatorname{sim}(q,d^{\gamma})}{\operatorname{sim}(q,d^{\varrho})+\operatorname{sim}(q,d^{\gamma})}$
Here, $\operatorname{sim}(q,d^{\varrho})$ is the similarity between generated context and question, and $\operatorname{sim}(q,d^{\gamma})$ is for retrieved context. Figure 8 illustrates the relationship between the average $\Delta\operatorname{sim}$ within each slice and the corresponding $\mathrm{DiffGR}$. From the results, we observe a general trend that on slices with a larger average similarity gap, LLMs exhibit a increased preference for generated contexts. Overall, the notable variations in $\mathrm{DiffGR}$ with respect to the similarity gap indicate that text similarity is a significant factor in the preference for generated contexts. These findings suggest that generated contexts should be applied with greater caution to mitigate the influence of highly relevant but misleading information.
To facilitate understanding why the similarity affects LLMs’ preference, we include some examples in Appendix A.11.4. From these cases, we observe that contexts with greater similarity often support candidate answers in a more straightforward manner, for instance, by mirroring the phrasing used in the questions. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences.
5.3 Effect of Context Completeness
In all above experiments, there is a key difference between generated and retrieved contexts that may affects the context preference: semantic and sentence completeness. Concretely, current retrieval systems typically employ fixed-length truncation to divide a complete article into multiple passages, which serve as the fundamental units for retrieval tasks (Karpukhin et al., 2020; Wang et al., 2019; Zhu et al., 2021). This truncation often results in retrieved contexts with incomplete semantic meaning, as well as sentences that are cut off at beginnings or endings. In contrast, generated contexts in above experiments are naturally produced by LLMs (Nature), resulting in enhanced semantic and sentence completeness.
Context Length Completeness Similarity Semantic Sentence Jaccard BERTScore Retrieved 107.4 0.114 0.801 Nature 109.7 0.202 0.879 Trunc 107.4 0.196 0.877 S-Trunc 105.9 0.193 0.876
Table 4: Contexts Statistics on NQ datasets. Appendix A.5 shows statistics on TQA, with similar conclusion.
Context Pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Nature vs. Ret 0.6562 0.5987 0.6101 0.5799 Trunc vs. Ret 0.2536 0.0958 0.3581 0.2042 S-Trunc vs. Ret 0.2394 0.1071 0.3302 0.1958
Table 5: $\mathrm{DiffGR}$ with different completeness in generated context. “Nature”, “Trunc.” and “S-Trunc.” represent three types of generated contexts with different completeness. “Ret” means retrieved contexts.
To investigate the potential effects of completeness on the observed bias, we conduct controlled experiments that vary the semantic and sentence completeness of generated contexts [^2] We also attempted to control the completeness of retrieved contexts, but it was challenging to isolate it from confounding factors like length. This aspect is left for future work. using following methods: (a) Truncation (Trunc.) eliminates the length constraints from the generation prompt of Section 3, allowing LLMs to generate extended contexts. These generated contexts are then truncated to match the length of retrieved contexts, thereby simulating both semantic and sentence incompleteness of retrieved contexts. (b) Sentence Truncation (S-Trunc.): Based on method (a), we truncate generated contexts only at the end of a sentence to preserve the sentence completeness, while simulating the semantic incompleteness.
Table 4 demonstrates that three types of generated contexts have similar average lengths and similarities. This suggests that the influences of similarity and length are mitigated, thereby highlighting the principal disparities in semantic content and sentence completeness.
We evaluate LLMs’ preference between generated versus retrieval context, varying the completeness of generated context, following the same pipeline in Section 4.1. Table 5 presents the $\mathrm{DiffGR}$ with different semantic and sentence completeness in generated contexts. A comparison between “Trunc.” and “S-Trunc.” reveals that sentence completeness does not significantly affect LLMs’ preference for generated contexts. In contrast, comparing “Nature” and “S-Trunc.”, we find a significant increase in preference for generated contexts, when they are semantically more complete. These findings indicate that LLMs are sensitive to the semantic completeness of the contexts, underscoring the necessity to investigate improved passage segmentation methods that maintain semantic completeness for current retrieval-augmented LMs.
6 Conclusion and Future work
In this study, we proposed a framework to investigate the underlying mechanisms by which LLMs merge retrieved and generated contexts. Our results reveal a pronounced bias towards generated contexts in several state-of-the-art LLMs (GPT 3.5/4, Llama2-7b/13b). We further identify two key factors contributing to this bias: higher similarity between generated contexts and questions, as well as the semantic incompleteness of retrieved contexts.
Our insights highlight the critical need for advanced integration methods that can validate and leverage information from both sources, moving beyond the current overreliance on generated contexts. Additionally, we find that LLMs display significant sensitivity to the semantic completeness of input contexts. This sensitivity necessitates improved passage segmentation strategies in current retrieval-augmented systems, thereby ensuring the preservation of intended meaning and the maximization of utility. Finally, addressing the challenges posed by highly relevant yet incorrect information generated by LLMs is an important direction for future research. It is crucial to develop methods for detecting and discounting misleading information produced by LLMs, especially as the volume of such content continues to escalate.
7 Limitation
This study investigates the context preference of LLMs when utilizing generated and retrieved contexts, with the aim of informing future work on retrieval-augmented and generation-augmented methods. Our work has the following limitations:
  • More Task: Our experiments and analyses have been conducted open-domain question answering, a classical knowledge-intensive task. It is unclear whether LLMs show similar preference in other tasks.
  • Solution: Since this study focus on analyzing the phenomena and causes of context preference, it does not propose specific solutions to effectively mitigate this preference. Future efforts will explore various methods to alleviate such problem.
  • More Complex Conflict Scenarios: As this work preliminarily focus on the conflicts between generated and retrieved contexts, we employ one instance for each types of contexts. When multiple contexts exist from a single source, conflicts may also arise among contexts from the same sources. How LLMs manage such complex conflicts need further investigation.
Appendix A Appendix
A.1 Related work
A.2 Generation-Augmented Approaches
Generation-augmented approaches prompt LLMs to generate intermediate contexts for the final response, thereby leveraging their extensive parametric knowledge acquired during the pre-training phase on vast text corpora (Roberts et al., 2020; Petroni et al., 2019). Recent studies have demonstrated the effectiveness of these methods across a range of tasks. Liu et al. (2022) represents an early exploration into augmenting LLMs in commonsense reasoning tasks using knowledge generated by LLMs themselves. Sun et al. (2023); Yu et al. (2022) employ LLMs to produce background documents (or recitations) and subsequently use these documents to enhance LLM performance in knowledge-intensive tasks. Another researches, known as the chain-of-thought (Wei et al., 2022; Kojima et al., 2022), prompts LLMs to generate intermediate reasoning steps to enhance LLMs’ reasoning abilities. Desipte the effectiveness of these methods, the LLM-generated knowledge may contain hallucinations (Ji et al., 2023) due to LLMs’ outdated memory (De Cao et al., 2021) and limited memory for long-tail knowledge (Kandpal et al., 2023). The LLM-generated inaccuracy information could potentially mislead current retrieval model (Dai et al., 2023) and open-domain question answering systems (Pan et al., 2023).
A.3 Retrieval-Augmented Approaches
The retrieval-augmented approaches (Guu et al., 2020; Lewis et al., 2020; Ram et al., 2023; Gao et al., 2023) enhance LLMs by incorporating relevant documents from external corpus. These approaches represent a promising direction for addressing the knowledge limitations of LLMs, such as the need for knowledge updates (Jang et al., 2022) and long-tail knowledge (Kandpal et al., 2023). Early retrieval-augmented methods (Guu et al., 2020; Lewis et al., 2020; Izacard et al., 2022) focused on the jointly training of LLMs and retrieval modules to improve their cooperation. With the evolution of general-purpose LLMs, recent studies (Ram et al., 2023; Shi et al., 2023) investigate the strategy of appending relevant documents directly to the input while keeping the LLMs frozen. Despite the effectiveness of these methods, the retrieval-augmented approaches still face challenges due to irrelevant retrieval results and incomplete knowledge coverage Yu et al. (2023); Mallen et al. (2023). These noisy retrieval results can misguide the outcomes of LLMs (Mallen et al., 2023; Yoran et al., 2023; Ren et al., 2023).
A.4 Hybrid Approaches and Knowledge Conflicts
Recent works investigate to merge retrieved and generated contexts to leverage both parametric knowledge and external knowledge Abdallah and Jatowt (2023); Yu et al. (2022). These combination methods have shown improved performance over those relying solely on a single information source in fully-supervised setting. Furthermore, (Zhang et al., 2023) proposes an improved method to effectively leverage the two sources of information, especially when conflicts arise. While these works have focused on improving the efficacy of hybrid approaches, the underlying mechanisms by which LLMs process conflicting information from different types of contexts remain underexplored.
Current research on knowledge conflicts in LLMs primarily focuses on two aspects: conflicts between input contexts and LLMs’ parametric memory, and conflicts among the input contexts themselves. Regarding the former, Xie et al. (2023); Chen et al. (2022) find that LLMs are highly receptive to the input contexts rather than their internal memory. Concerning conflicts within multiple input contexts, Chen et al. (2022) demonstrates that LLMs tend to rely on a few most relevant retrieved contexts. Additionally, Xie et al. (2023) reveals a confirmation bias, i.e., LLMs demonstrate a tendency to favor contexts that align with their parametric knowledge, when confronted with both supporting and opposing contexts. However, these studies are limited to analyzing context conflicts within a single type of input contexts. Our work complements these studies by considering conflicts between generated and retrieved contexts, and reveals several key factors, such as semantic completeness and text similarity.
A.5 Length Control for Generated Contexts
A.5.1 Length Distribution across LLMs
Dataset Retrieved Generated GPT 4 GPT 3.5 Llama2-13b Llama2-7b NQ 107.3 108.0 106.0 110.1 104.0 TQA 106.3 107.2 104.9 105.5 102.6
Table 6: Average lengths of the generated and retrieved contexts. Length is measured in the number of words after punctuation removal.
In our proposed framework, we regulate the length of generated contexts by incorporating length constraint in the prompt: Generate a background context from Wikipedia to answer the given question {#question} Keep the length of the document around $n$ words We observed that GPT 4 effectively controls the output length, whereas other models struggle with this aspect. To address this issue in the latter, we employ multiple values of $n$ and select the one that best matched the retrieved context.
As the result, Figure 9 shows the length distribution of retrieved contexts and contexts generated by various LLMs. The length distribution of retrieved contexts is more concentrated due to the fact that these contexts consist of text truncated precisely at one hundred words, coupled with their titles (Karpukhin et al., 2020). Variations in the length of different retrieved articles are solely attributable to the differences in title lengths.
A.6 Length Distribution with Different Control Methods
Figure 10 illustrates the length distribution for generated contexts corresponding to three different length control methods. From the results, we can observe that the contexts generated by GenRead (Yu et al., 2022) are significantly longer compared to the retrieved contexts. This difference in length could potentially influence the model’s bias or inclination.
Table 7 illustrates the statistics of three types of generated contexts with different methods to control the length. These contexts show similar average length and similarity, with the preliminary difference in the completeness.x
Context Completeness NQ-AIR TQA-AIR Sentence Semantic Length Jaccard Length Jaccard GPT-3.5 Nature 106.0 0.187 105.0 0.398 S-Trunc 104.9 0.184 106.3 0.397 Trunc 107.4 0.187 103.6 0.397 Llama2-13b Nature 109.7 0.202 105.5 0.313 S-Trunc 105.9 0.193 104.8 0.2934 Trunc 107.4 0.196 106.3 0.2954
Table 7: Average length and similarity of generated context with different control methods. Three types of methods create context with comparable average length and similarity. The core difference lies in the completeness.
Figure 9: Length distribution of generated and retrieved contexts on different datasets with different generator models.
Figure 10: Length distribution of generated and retrieved contexts on the NQ dataset with GPT-3.5 as the generator. "genread" represents the contexts generated by the original GenRead method (Yu et al., 2022). "trunc", "strunc", and "nature" are the generated contexts using three different methods to control the length.
A.7 Dataset Statistics
Table 8 presents the data size of context-conflicting datasets corresponding to various generator-reader pairs. The statistics indicate that conflicting data comprise a substantial proportion across all combinations of generators and readers.
Reader Generator NQ (3610) TQA (11313) NQ-AIG NQ-AIR TQA-AIG TQA-AIR GPT 4 GPT 4 1120 763 1712 681 GPT 4 GPT 3.5 1017 922 - - GPT 4 Llama2-13b 730 1461 - - GPT 4 Llama2-7b 600 1627 - - GPT 3.5 GPT 4 1514 769 2701 794 GPT 3.5 GPT 3.5 1337 857 2389 1042 GPT 3.5 Llama2-13b 875 1318 1781 2119 GPT 3.5 Llama2-7b 701 1502 1471 2641 Llama2-13b GPT 4 2501 767 4769 741 Llama2-13b GPT 3.5 2211 899 4210 1038 Llama2-13b Llama2-13b 1441 1336 2982 2091 Llama2-13b Llama2-7b 1221 1583 2567 2773 Llama2-7b GPT 4 2699 668 5370 830 Llama2-7b GPT 3.5 2435 785 4813 1120 Llama2-7b Llama2-13b 1569 1220 3526 2051 Llama2-7b Llama2-7b 1423 1381 3064 2604
Table 8: The data quantities of the constructed subsets for different (Generator, Reader) pairs. NQ and TQA refer to the original complete test sets.
A.8 Effect of Context Order
Order NQ-AIR TQA-AIR
generated-first 69.9 68.2
retrieved-first 66.5 55.6
random 69.1 58.6
Table 9: The generation ratio (%) with different context order on NQ-AIR and TQA-AIR datasets. The backbone LLM is GPT 3.5.
In above experiments, retrieved and generated contexts are presented with random order. Previous studies Xie et al. (2023); Liu et al. (2023); Lu et al. (2022) have found that the model may be sensitive to the order of the input contexts. In their experiments, the input context was either all retrieved Liu et al. (2023) or all generated Xie et al. (2023). We conducted experiments to investigate whether the context order impacts the preference for the generated context. The generated and retrieved contexts are concatenated with three different orders: generated-first, retrieved-first and random. We compute the context preference with different context order respectively.
As shown in Table 9, across all context orders, LLMs consistently show a strong tendency to generated contexts (generation ratio > 0). When the retrieved context is positioned first, there is a slight reduction in the generation ratio. This reduction may result from the LLMs’ preference for generated contexts being partially offset by their bias towards the top context (Liu et al., 2023; Xie et al., 2023).
A.9 More Results on AIG Datasets
Figure 11 shows the $\mathrm{DiffGR}$ with different (reader, generator) pairs on their corresponding NQ-AIG datasets. It can be observed that LLMs show strong tendency to rely on generated contexts across various (reader, generator) pairs.
Figure 11: $\mathrm{DiffGR}$ (%) with different (reader, generator) pairs on their corresponding NQ-AIG datasets.
A.10 More Details about Confirmation Bias
A.10.1 Experiment Setting
We construct counter-memory answers and contexts based on the original AIR context-conflicting subsets, as outlined below:
Counter-Memory Answers Preparation. For each question on the AIR subsets, we substitute the original memory answer (e.g., “Canada”) with a same-type yet distinct entity (e.g., “United States”), which serves as the counter-memory answer. Concretely, we employ ChatGPT to associate a different entity of the same type. It is noteworthy that the prompt does not incorporate the question, ensuring that the LLM’s counter-memory answer is merely a categorical equivalent rather than a memory related to the question.
Give you a reference word, transform it into a analogous word. Reference: Missouri River Analogous Word: Mississippi River Reference: {memory answer} Analogous Word:
Furthermore, to facilitate the calculation of $\mathrm{DiffGR}$, we only retain those samples where the counter-memory answer also diverged from $a_{\phi}^{\gamma}$, the answer provided by the retrieved context. Formally, the counter-memory answer satisfies $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\varrho}$ and $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\gamma}$.
Counter-Memory Contexts Generation. We enforce LLMs to generate a counter-memory context that supports the counter-memory answer. This approach is inspired by the success of using LLMs to generate misinformation (Pan et al., 2023).
Generate a background document in support of the given opinion to the question. Keep the length of the document around $n$ words. Question: {question} Opinion: {counter memory answer} Document:
Following Section 3, the prompt also incorporates a length constraint to mitigate the influence of length, as evidenced by the length statistics presented in Table 10.
Answer Consistency Checking. To verify that the counter-memory context actually supports the counter-memory answer, we retain only those instances where the predicted answer, derived exclusively from the counter-memory context, exactly match the counter-memory answer:
$a_{\phi}^{\varrho^{\prime}}=\phi(q,d_{1}^{\varrho^{\prime}})$
Context NQ-AIR TQA-AIR Length Jaccard Length Jaccard Retrieved 107.5 0.111 106.3 0.184 GPT-3.5 Gen 105.9 0.194 104.9 0.385 Ctr 105.7 0.251 103.8 0.499 Llama2-13b Gen 107.4 0.197 104.3 0.319 Ctr 106.7 0.165 103.5 0.310
Table 10: Average length and Jaccard similarity of different contexts. Gen, Ret and Ctr respectively represent generated contexts, retrieved contexts and counter-memory contexts. Detailed discussion about Jaccard similarity is shown in Section 5.2.
Reader & Generator NQ-AIR TQA-AIR
GPT-3.5 504 596
Llama2-13b 883 1376
Table 11: The size of context-conflicting subsets with counter-memory contexts.
Following the above procedures, we obtain a subsets from original AIR datasets, with each instance encompassing a question $q$, a generated context $d_{1}^{\varrho}$, a retrieved context $d_{1}^{\gamma}$, a counter-memory context $d_{1}^{\varrho^{\prime}}$ and the associated answers for these contexts.
A.10.2 Discussion
In Table 3, we observe another phenomenon: GPT-3.5 exhibits a stronger inclination towards counter-memory contexts compared to the original generated contexts, which does not exist in Llama2-13b. This phenomenon is also found in the results of Xie et al. (2023). Upon further investigation, we discovered that this may be due to the relatively higher similarity betwee the counter-memory context generated by GPT-3.5 and the question, as shown in Table 4. We also analyse the conflictQA datasets proposed by Xie et al. (2023), and also find the similarity gap between generated and counter-memory contexts. This means that previous works may ignore some significant factors to LLMs’ preference, such as similarity (investigated in Section 5.2).
A.11 More Details and Results about Similarity
A.11.1 Similarity Metric
We employ Jaccard similarity to assess the term-based overlap, and BERTScore (Zhang et al., 2020) for evaluating the semantic similarity between contexts and questions. To mitigate the effect of length discrepancies between contexts and questions, we calculate the similarity at the sentence level and then aggregate them to derive the overall context-question similarity. In this section, we adopt a maximum aggregation strategy due to the single-hop nature of the NQ and TQA datasets, where the majority of questions can be answered using a small subset of sentences. We also try average aggregation strategy and observe similar results.
Figure 12 illustrates the distribution of similarity when employing maximum and average aggregation methods. It is observable that the generated contexts exhibit a markedly higher degree of similarity regardless of the aggregation method used. Furthermore, this disparity in similarity is more pronounced with maximum aggregation, as contexts typically contain sentences that are irrelevant, which dilute the similarity scores when an average aggregation is applied.
Figure 12: Similarity distribution with maximum or mean aggregation strategies. Generated contexts consistently exhibit higher similarity across two aggregation strategies.
A.11.2 Similarity Distribution
Figure 13 and 14 show the similarity distribution of retrieved and generated contexts across various generators. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
A.11.3 Effect of Similarity.
Figure 15 demonstrates a general trend that on slices with a smaller average similarity gap, LLMs exhibit a reduced preference for generated context. This trend is not very stable on NQ-AIR, which can be attributed to the smaller number of instances on NQ-AIR, as shown in Table 8.
Figure 13: BERTScore distribution of retrieved context and context generated by different LLMs. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
Figure 14: Jaccard Similarity distribution of retrieved context and context generated by different LLMs. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
(a) GPT 4
(b) GPT 3.5
(c) Llama2-13b
(a) GPT 4
(b) GPT 3.5
(c) Llama2-13b
(d) Llama2-7b
Figure 15: The generation ratio in slices with different average $\Delta\operatorname{sim}$. $\Delta\operatorname{sim}$ is the difference in similarity between the generated context and the retrieved context.
A.11.4 Cases about Similarity
Table 12 shows examples that contain contexts with different similarity to the question. The contexts with high similarity typically directly support answering by repeating the phrasing in the question. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences. These observations indicate that text similarity can partly reflect the relevance between a question and a context, as well as the difficulty a LLM encounters in identifying potential answers.
TQA-AIR Example TQA-AIG Example Question Between 1959 and 1967 which city was the capital of Pakistan (Islamabad was being built)? Who is the most successful UK solo artist in the USA? Golden Answer Rawalpindi Elton John Generated Context Between 1959 and 1967, the capital of Pakistan was Karachi. Karachi is the largest city in Pakistan and is located on the southern coast of the country… Jaccard Similarity: 0.47 BertScore: 0.93 Elton John is the most successful UK solo artist in the USA. Born Reginald Kenneth Dwight in 1947, he adopted the stage name Elton John in the late 1960s… Jaccard Similarity: 0.69 BertScore: 0.93 Retrieved Context …was first shifted temporarily to Rawalpindi in the early 60s, and then to Islamabad when essential development work was completed in 1966… Jaccard Similarity: 0.16 BertScore: 0.85 …In 2009, Jay Sean’s single "Down" reached the number one spot on the "Billboard" Hot 100 and sold millions in the United States, making him ẗhe most successful male UK urban artist in US chart historyät the time… Jaccard Similarity: 0.14 BertScore: 0.86 Model output Karachi Elton John
Table 12: Some examples where both the generator and reader are GPT-3.5. We highlight the incorrect candidate answers in the context in pink, and the correct answers in the context in green.